# Fatigue-PINN: Physics-Informed Fatigue-Driven Motion Modulation and Synthesis

### v1

@@ V-A Ablations
Within Fatigue-PINN, we modeled our ID and FD networks as PINNs as well. The primary distinction in our corresponding BiLSTM networks lies in the loss function, where, in addition to the MSE term, which minimizes the loss between the neural network estimations and the real values of joint actuator torques (for ID) and joint angles (for FD), we incorporated a physics-based loss term derived from the classical equation of motion (Equation (15)). This equation denotes the relationship of joint kinematics (generalized positions, i.e. joint angles, velocities, and accelerations), with generalized forces (i.e. joint torques) during human motion:
@@
Nevertheless, the changes in model accuracy and generated fatigued motion are insignificant whether we use PINNs to formulate our ID and FD models or not (the NRMSE values of the ID/FD PINN models differ by about ±$\pm 5\%$ compared to those in Table I). Therefore, even though we utilized the standard BiLSTM architecture for the ID and FD models, their PINN versions offer a viable solution in cases of limited data availability. This is because PINN loss functions act as soft physical constraints during human motion, facilitating the inference of corresponding joint torques and angles.
It is worth mentioning that we also experimented with training the 3CC-$\lambda$ unsupervisingly (solving the forward problem), where no ground-truth data for fatigued, $M_{F}$, and resting state $M_{R}$, are provided. In this case, the PINN loss, $L$ (see Equation (9)), still has its physics-based term, $L_{PB}$ (Equation (11)), and instead of $L_{NN}$ (Equation (10)), we add a loss term to incorporate boundary conditions regarding $M_{F}$ and $M_{R}$. However, training 3CC-$\lambda$ unsupervisinlgy requires more computational resources compared to the approach outlined in Section III-B.
@@ V-B Fatigued Movements
Fatigue is usually neglected in most musculoskeletal models as well as in data-driven motion synthesis frameworks, leading in bypassing an element that could significantly boost animation realism. The widely emerging Physics-Informed Neural Networks are also making their debut in human biomechanics estimation tasks (e.g. prediction of joint angles/forces/moments, muscle forces, etc. [12, 13, 14, 52, 53, 54, 55, 56]) since they can inject physics-based domain knowledge into the learning procedure to both approximate PDEs describing the problem-to-solve (e.g. Inverse of Forward Kinematics equations) and penalize solutions not adhering to physical laws, limiting the solution space, which leads to faster convergence during training. Consequently, PINNs provide more accurate physically consistent solutions, while requiring a significantly smaller amount of data to approximate the underlying physical laws.
@@
In this study, we present Fatigue-PINN, an advanced end-to-end PINN-based deep learning framework to synthesize fatigued human motion without relying on fatigued motion capture data for training. Our architecture is composed of three modules, the Inverse Dynamics Module, which translates input joint angles into joint torques, the Fatigue module, i.e. the 3CC-$\lambda$ model to simulate the effects of fatigue in the maximum exerted joint torques, and the Forward Dynamics model, which transforms these fatigued joint torques back into joint angles, thus enabling the fatigued animation of a 3D character. 3CC-$\lambda$ is a PINN adaptation of the 3CC state machine [22, 23, 24], augmented with a method to account for joint-specific fatigue configurations, thereby ensuring seamless incorporation of fatigue effects in different motions and avoid sharp step-wise motions. Additionally, we use ID and FD models employed upon each DoF. These joint-specific models enhance predictive accuracy and are also compatible with the joint-specific fatigue profiles as defined by 3CC-$\lambda$. By having control at the joint level, our model is capable of generating physically plausible movements with reduced motion artifacts. The ID and FD models are LSTM-based to extract the temporally evolving fatigue impact and can be utilized as surrogate models to perform the respective tasks without the need of GRFs.
@@
Our results indicate that Fatigue-PINN produces realistic fatigue effects on open-type movements, which are consistent with the existing literature [70, 18, 19, 3, 2]. It is worth mentioning that in our experiments, all upper body muscles were fatigued isometrically, causing overlapping kinematic changes to render the externally perceived fatigue impact more evident. Nevertheless, since our Fatigue-PINN framework is joint-specific it can also be utilized to study the effects of fatigue on posture and joint kinematics after inducing localized muscular fatigue (e.g. fatiguing shoulder and observing the impact on elbow and trunk), just as in [2], where no data-driven method was used.
@@
Furthermore, our end-to-end Fatigue-PINN framework can be utilized to both facilitate the work of animators and aid biomedical engineers with investigating the effects of externally perceived fatigue on posture and movements. By simply providing a motion sequence (i.e. a sequence of joint angles), the fatigued movement at a specified level of fatigue, is synthesized. Fatigue-PINN does not require direct input of joint torques for fatigue incorporation, as it can convert joint angles into torques and vice versa. The latter enables the seamless integration into motion synthesis pipelines that operate on joint angles.
@@
A limitation of our work is that even though it succeeds in modeling fatigue in closed-type movements, i.e. when an object is involved, such as lifting or pushing a box, it lacks in accuracy, with motion artifacts being present, mostly in the end effectors. A possible solution to fix hand artifacts is to incorporate in our framework contact points [27, 28] as well as forces produced by the human-object interactions. As for foot contact artifacts, exploiting GRFs and foot contact information [32] as well as utilizing GANs [72] or diffusion models [37], are possible ways to mitigate these artifacts. The latter is proposed as future work.
@@ VI Conclusion
Addressing the gap of data-driven methods for modeling fatigue in literature, we introduce Fatigue-PINN, a PINN-based deep learning framework for synthesizing fatigued human movements without explicitly modeling the complex underlying physics laws governing the human body. Our framework consists of three modules, namely Inverse Dynamics, Fatigue, and Forward Dynamics modules, forming an encoder-decorder-like architecture for producing fatigued joint angle sequences based on fatigued-induced joint torques. Despite the incorporation of fatigue within the torque space, the end-to-end encoder-decoder configuration processes joint angles to produce their fatigued adaptations, facilitating the integration with any motion synthesis model functioning on joint angles and testing the real-time capabilities of the system. To estimate the fatigued state of motion, we model the 3CC state machine as a PINN, 3CC-$\lambda$, and we enhance it with the parameter $\lambda$ to characterize the fatigue profile of each joint and contribute to the smooth integration of fatigue in produced motion, thus, preventing abrupt frame transitions. Our semi-dynamics joint-specific approach, employing an ID and FD BiLSTM model for each joint, not only improves the overall predictive accuracy but also enables better alignment with joint-specific fatigue configurations and handling motion artifacts. Fatigue-PINN simulates the effects of externally perceived fatigue on open-type human movements comparable with findings from real-world experimental fatigue studies, even though it does not require fatigued motion capture data for the learning process. The adaptation of Fatigue-PINN to closed-type movements involving contacts and forces in human-object interactions remains as a future research direction.
Acknowledgments
The data used in this project was obtained from https://fling.seas.upenn.edu/~mocap/cgi-bin/Database. The research project is implemented in the framework of H.F.R.I call “Basic research Financing (Horizontal support of all Sciences)” under the National Recovery and Resilience Plan “Greece 2.0” funded by the European Union – NextGenerationEU (H.F.R.I. Project Number: 16469.).

### v2

@@ V-A Ablations
Within Fatigue-PINN, we modeled our ID and FD networks as PINNs as well. The primary distinction in our corresponding BiLSTM networks lies in the loss function, where, in addition to the MSE term, which minimizes the loss between the neural network estimations and the real values of joint actuator torques (for ID) and joint angles (for FD), we incorporated a physics-based loss term derived from the classical equation of motion (Eq. (15)). This equation denotes the relationship of joint kinematics (generalized positions, i.e. joint angles, velocities, and accelerations), with generalized forces (i.e. joint torques) during human motion:
@@
Nevertheless, the changes in model accuracy and generated fatigued motion are insignificant whether we use PINNs to formulate our ID and FD models or not (the NRMSE values of the ID/FD PINN models differ by about $\pm 5\%$ compared to those in Table IV). Therefore, even though we utilized the standard BiLSTM architecture for the ID and FD models, their PINN versions offer a viable solution in cases of few training data. This is because PINN loss functions act as soft physical constraints during human motion, facilitating the inference of corresponding joint torques and angles.
It is worth mentioning that we also experimented with training the 3CC-$\lambda$ unsupervisingly (solving the forward problem), where no ground-truth data for fatigued, $M_{F}$, and resting state $M_{R}$, are provided. In this case, the PINN loss, $L$ (see Eq. (9)), still has its physics-based term, $L_{PB}$ (Eq. (11)), and instead of $L_{NN}$ (Eq. (10)), we add a loss term to incorporate boundary conditions regarding $M_{F}$ and $M_{R}$. To evaluate which of the two training schemes, supervised or unsupervised, is more efficient for the 3CC-$\lambda$ model, we calculated the Pearson’s Correlation Coefficient, $R^{2}$, between predicted and ground-truth (experimental) values. In the unsupervised scheme, the $R^{2}$ for each joint angle across emotional trials of an action (punching, waving, or throwing) range from $\sim 0.93-0.96$, while the supervised version of the model has $R^{2}$ values of $\sim 0.91-0.97$. Hence, the two approaches provide 3CC-$\lambda$ models whose $R^{2}$ values differ by $\sim 0.01-0.03$ units. This small difference in accuracy between the two 3CC-$\lambda$ versions is probably due to the fact that the equations driving the 3CC are relatively simple, facilitating the convergence to the correct solution, even in the unsupervised version. The latter enables us to follow either of the two training processes for our 3CC-$\lambda$ model, however, training 3CC-$\lambda$ unsupervisingly requires more computational resources compared to the approach outlined in Section III-B.
@@ V-B Fatigued Movements
Fatigue is usually neglected in most musculoskeletal models as well as in data-driven motion synthesis frameworks, leading in bypassing an element that could significantly boost animation realism. The widely emerging Physics-Informed Neural Networks are also making their debut in human biomechanics estimation tasks (e.g. prediction of joint angles/forces/moments, muscle forces, etc. [12, 13, 14, 52, 53, 54, 55, 56]) since they can provide more accurate physically consistent solutions for both forward and inverse problems, while requiring a significantly smaller amount of data to approximate the underlying physical laws.
@@
In this study, we present Fatigue-PINN, an advanced end-to-end PINN-based deep learning framework to synthesize fatigued human motion without relying on fatigued motion capture data for training. Our architecture is composed of three modules, the Inverse Dynamics Module, which translates input joint angles into joint torques, the Fatigue module, i.e. the 3CC-$\lambda$ model to simulate the effects of fatigue in the maximum exerted joint torques, and the Forward Dynamics model, which transforms these fatigued joint torques back into joint angles, thus enabling the fatigued animation of a 3D character. 3CC-$\lambda$ is a PINN adaptation of the 3CC state machine [22, 23, 24], augmented with a method to account for joint-specific fatigue configurations, thereby ensuring seamless incorporation of fatigue effects in different motions and avoid sharp step-wise motions. Additionally, we use ID and FD models employed upon each DoF. These joint-specific models enhance predictive accuracy and are also compatible with the joint-specific fatigue profiles as defined by 3CC-$\lambda$. By having control at the joint level, our model is capable of generating physically plausible movements with reduced motion artifacts. The ID and FD models are LSTM-based to extract the temporally evolving fatigue impact and can be utilized as surrogate models to perform the respective tasks without the need for GRFs. GRFs are expected to vary between non-fatigued movements and fatigued movements, to account for changes in the body’s center of mass/pressure, due to variations in posture and kinematic patterns [70, 76]. In simpler words, alterations in upper limb torques (the ones investigated in this work) due to fatigue, will influence GRFs. Thus, in order to include GRFs in the training/inference of our ID and FD models, the use of training datasets that contain both fatigued motion capture and GRF data, is required. Such datasets with motion sequences covering open-type movements, as the ones explored in this work, are not available in the literature.
@@
Our results indicate that Fatigue-PINN produces fatigue effects on open-type movements quite similar to those of externally perceived fatigue, which are consistent with the existing literature [70, 18, 19, 3, 2]. It is worth mentioning that in our experiments, at each action (i.e. waving, punching, and throwing), all upper body muscles were fatigued isometrically, causing overlapping kinematic changes to render the externally perceived fatigue impact more evident. For instance, for the punching motion, the $\lambda$ factor for all upper body joints is set to $0.6$, which provides a good balance between motion smoothness and conveyance of externally perceived fatigue. Nevertheless, since our Fatigue-PINN framework is joint-specific it can also be utilized to study the effects of fatigue on posture and joint kinematics after inducing localized muscular fatigue (e.g. fatiguing the shoulder and observing the impact on elbow and trunk), just as in [2], where no data-driven method was used.
A perceptual study was also conducted indicating that the participants were able to distinguish between the various levels of fatigue of a motion and further characterize which fatigued motion exhibited the most pronounced fatigue effects. Moreover, the participants succeeded in specifying the latter effects in fatigued motion sequences across punching, throwing, and waving motion types, as well as providing their own visual observations, pinpointing fatigued effects that the authors were incapable of defining. However, it is worth mentioning that fatigued throwing movements, posed a challenge to the responders, rendering them unable to achieve high percentages of accurate distinctions between fatigued and non-fatigued throwing motions. Despite this fact, the majority of participants described the fatigued motion sequences of all motion types (punching, throwing, and waving) as “realistic”. Moreover, we assume that providing the participants with longer-in-duration animated videos, at different frame speeds, as well as featuring multiple view angles of the same motion, would aid the responders to perceive more fatigued-induced elements in the respective motions. The outcome of the perceptual study was further supported by a statistical analysis based on $\chi^{2}$ and One-Way ANOVA tests.
@@
Furthermore, our end-to-end Fatigue-PINN framework can be utilized to both facilitate the work of animators and aid biomedical engineers with investigating the effects of externally perceived fatigue on posture and movements. By simply providing a motion sequence (i.e. a sequence of joint angles), the fatigued movement at a specified level of fatigue is synthesized. Fatigue-PINN does not require direct input of joint torques for fatigue incorporation, as it can convert joint angles into torques and vice versa. The latter enables the seamless integration into motion synthesis pipelines that operate on joint angles.
@@
A limitation of our work is that even though the prediction metrics of the Fatigue-PINN framework indicate a good accuracy in closed-type movements, i.e. object manipulation or multi-contact motions, this is not reflected in the animation results of such motions, with motion artifacts being present, mostly in the end effectors. More specifically, we evaluated the Fatigue-PINN framework on the ”lifting a box” and ”pushing a box” motion types from the Adapt Emotional Actions dataset. In both movements, the overall testing accuracy of Fatigue-PINN is quite high, around $92-95\%$, while the average -across motion trials- NRMSE and $R^{2}$ values of ID/FD models for right upper limb joint torques/angles are very close to those reported in Tables IV & V. As for the produced fatigue animation, since our framework does not support contact integration, in both box manipulation motions, the hands are misplaced, resulting in either penetrating or not touching the “ghost” of the box. Also, foot skating artifacts are present. A possible solution to fix hand artifacts is to incorporate in our framework contact points [27, 28] as well as forces produced by human-object interactions. As for foot contact artifacts, exploiting GRFs and foot contact information [32] as well as utilizing GANs [77] or diffusion models [37], are possible ways to mitigate these artifacts. The latter is proposed as future work.
Another limitation is that the dataset utilized in this work is recorded upon the expressive movements of one subject. Thus, even though the diverse motion trials improve the generalization ability of our ID and FD models across variations of the same motion (i.e. the same movement expressed via different emotions), the absence of subject variability hinders the evaluation of our model in terms of generalization across subjects. In addition, a dataset containing more motion trials for each action movement would be preferable to further enhance the robustness of the model. More experiments on larger datasets with both subject and motion trial diversity remain as future work.
Finally, the Adapt Emotional Actions dataset does not provide any fatigued motion capture data, and, as far as the authors are concerned, fatigued motion capture datasets covering open-type movements as the ones utilized in this work, are not present in the literature or open-source. The inability to validate our framework against real fatigued motion capture data is also a limitation of our work, and we are currently working on an expansion of Fatigue-PINN, incorporating fatigue in the lower limbs, such as gait movements, where actual fatigued data are available through numerous publications. In the latter case, the comparison between our framework and baseline methods will be facilitated.
@@ VI Conclusion
Addressing the gap of data-driven methods for modeling fatigue in literature, we introduce Fatigue-PINN, a PINN-based deep learning framework for synthesizing fatigued human movements without explicitly modeling the complex underlying physics laws governing the human body. Our framework consists of three modules, namely Inverse Dynamics, Fatigue, and Forward Dynamics modules, forming an encoder-decoder-like architecture for producing fatigued joint angle sequences based on fatigued-induced joint torques. Despite the incorporation of fatigue within the torque space, the end-to-end encoder-decoder configuration processes joint angles to produce their fatigued adaptations, facilitating the integration with any motion synthesis model functioning on joint angles and testing the real-time capabilities of the system. To estimate the fatigued state of motion, we model the 3CC state machine as a PINN, 3CC-$\lambda$, and we enhance it with the parameter $\lambda$ to characterize the fatigue profile of each joint and contribute to the smooth integration of fatigue in produced motion, thus, preventing abrupt frame transitions. Our semi-dynamics joint-specific approach, employing an ID and FD BiLSTM model for each joint, not only improves the overall predictive accuracy but also enables better alignment with joint-specific fatigue configurations and handling motion artifacts. Fatigue-PINN simulates the effects of externally perceived fatigue on open-type human movements comparable with findings from real-world experimental fatigue studies, even though it does not require fatigued motion capture data for the learning process. The adaptation of Fatigue-PINN to closed-type movements involving contacts and forces in human-object interactions as well as periodical movements where lower limbs are mostly involved, like gait, remains a future research direction.
Appendix A Pipeline Details
The aim of this Appendix is to provide more insights regarding the procedure we followed to develop and train our Fatigue-PINN framework, in order to aid researchers with reproducing our results. The full Fatigue-PINN pipeline consists of the following modules:
  1. 1.
    Raw Data Pre-Processing
  2. 2.
    Musculoskeletal Modeling
  3. 3.
    Training Data Pre-Processing
  4. 4.
    Deep Learning Models’ Training
  5. 5.
    Animation in Unity Environment
as illustrated in Fig. 10.
Figure 10: An overview of our pipeline to develop and test the Fatigue-PINN framework.
A-A Raw Data Curation
The first step of our pipeline is to pre-process the raw motion capture and GRF data contained in the Adapt Emotional Actions open dataset, utilized for the purposes of this work. Since these data are in C3D format, we created a Python script that converts C3D files to motion file formats that the OpenSim musculoskeletal modeling software can process. In particular, these scripts make use of OpenSim Python API’s built-in functions to read C3D files and write marker data in TRC files and their respective force plate data in MOT files. By also using functions from the OpenSim Python library, marker, and GRF data are rotated over 90-90 degrees along the x-axis and $270$ degrees over the y-axis to match the dataset’s experimental coordinate system to the one of OpenSim. The raw data of each trial of every motion type investigated in this work is saved in individual motion files.
Before saving, GRF data are also converted from millimeters to meters to be aligned with the measurement units of OpenSim and smoothed via a Locally Weighted Regression algorithm. This method, also known as Locally Weighted Scatterplot Smoothing (LOWESS), fits a non-parametric regression curve to a scattered set of data. More specifically, unlike traditional models that apply a single global model, LOWESS fits multiple models over local regions of data, weighing nearby points higher based on their proximity to the target value (the value where we want to estimate the output). Thus, LOWESS flexibly adapts to complex datasets, handling noisy and non-linear data. In our implementation of LOWESS, we utilized a bell-shaped kernel function with parameter $c$, where larger $c$ values will result in a smoother curve. The aforementioned process, i.e. measurement unit conversion and smoothing, is implemented through Python scripting.
A-B Musculoskeletal Modeling Procedure
We used the OpenSim [45] open-source tool to derive the joint kinematics and dynamics (especially joint angles and torques), based on the markers’ spatial trajectories and GRFs that were processed through the above procedure (see Section A-A). As depicted in Fig. 10, the musculoskeletal modeling procedure includes the following analyses:
  • Scaling: Motion capture data containing the subject’s neutral pose marker data are used to scale a generic musculoskeletal model (that is, a Hill-type full-body presented in [60]), to the anthropometric measurements of the respective subject (e.g., length of each body segment, height, etc.). Hence, obtain a personalized model of the subject whose motion was recorded for the Adapt Emotional Actions dataset. In order to scale the dimensions of each body segment, the Scaling tool ensures that the distances between the virtual markers of the OpenSim model match the distances between the experimental markers (the ones from motion capture). To optimize the scaling process, the virtual markers’ positions were manually modified before applying the scaling tool. The output of this step is the scaled model.
  • Inverse Kinematics: The IK tool is employed upon the scaled model to reproduce the recorded motion from the experimental markers’ spatial trajectories (motion capture data). In particular, at each time frame, the error between experimental (motion capture) markers and musculoskeletal model markers’ position is calculated by IK to estimate the best matching pose. The latter process is carried out via a Weighted Least Squares equation, where a sum of marker weighted squared errors is minimized. Through the IK process, the generalized positions, i.e., joint angles of the whole body are extracted and stored in a MOT file format.
  • Inverse Dynamics: The joint angles computed via IK and the experimentally measured GRFs are used as an input to the Inverse Dynamics tool to calculate the generalized forces, i.e. joint torques, as mandated by Eq. (15). Joint torques are stored in a motion file of STO format.
The inputs and parameters for each analysis (e.g. input files, duration of the motion to be analyzed, etc.) are specified through XML files, automating the OpenSim pipeline.
A-C Training Data Pre-Processing
Following the musculoskeletal modeling process, joint kinematics and torques are extracted from experimental motion capture and force plate data. Before using this derivate data for the training of our ID/FD BiLSTM models, further processing is required. The processes outlined in this section are conducted using Python scripting since the models that constitute our Fatigue-PINN framework were trained in Python Keras.
The motion files containing generalized positions (MOT files) and forces (STO files) are converted to XLS files, which can be easily manipulated and converted to Dataframe format using the Python Data Analysis Library (Pandas) [78]. Joint angles and torques were normalized in the range $[0,1]$ using min-max normalization as indicated by the following mathematical formula (Eq. 16):
$x_{normalized}=\frac{x–\min{\{x_{1},\dots x_{T}\}}}{\max{\{x_{1},\dots x_{T}\}}-\min{\{x_{1},\dots x_{T}\}}}$ (16)
where $x$ is a sequence of either joint angle data $\mathbf{q}^{N}=\left\{\mathbf{q}^{N}_{1},...,\mathbf{q}^{N}_{T}\right\}$ or a sequence of joint torques $\mathbf{\tau}^{N}=\left\{\mathbf{\tau}^{N}_{1},...,\mathbf{\tau}^{N}_{T}\right\}$ for t=1,,Tt=1,...,T and $N=32$. Also, minmin is the minimum and maxmax is the maximum value of the joint angles/torques sequences, respectively.
As mentioned in Section III-A, the Adapt Emotional Action dataset provides multiple trials for each motion, which have different durations, i.e. number of frames, ranging from $\sim 450-600$ frames. Thus, for every motion type, e.g. punching, we resample all motion trials to have the same number of frames using linear interpolation based on the trial with the longest duration. Our linear interpolation algorithm was developed using built-in functions from Python’s SciPy Library [79]. The resampled motion trials will have a duration of approximately $600$ frames.
A-D Training Scheme for Deep Learning Models
[…]
The ID and FD models were evaluated through a three-fold cross-validation approach applied across all motion trials of each motion type. Explanatory, the trials of each action are split into $3$ folds where the $\frac{2}{3}$ is being utilized as training data and the left-out $\frac{1}{3}$ as test data. The process is iteratively executed until each subset has been utilized as test data. Subsequently, the prediction accuracy of our model is computed by averaging the results across the three validation folds. Mean Square Error was utilized as the loss function of our ID and FD models.
The 3CC-$\lambda$ was also trained supervisingly, utilizing datasets cited in [62], containing fatigued joint torques measured after fatiguing experimental protocols. These datasets have very few time frames, and they are not subjected to the above pre-processing procedure, yet they only require normalization in the range of $[0,1]$. Equations (9), (10), and (11) describe the PINN-loss of our 3CC-$\lambda$ model.
We trained our models with a batch size of $32$, for $1000$ epochs with early stopping, meaning that this number of epochs was never reached, avoiding over-fitting. Adam [65] was used as the optimizer with a learning rate of $0.001$.
Given that our 3CC-$\lambda$ model requires training based on datasets other than the Adapt Emotional dataset, it can be used as a pre-trained model. Therefore, the Fatigue-PINN framework can be either trained end-to-end or separately for each module.
A-E Animation in Unity Environment
To simulate our results in Unity Game Engine Environment [69], we may either:
  • Re-implement our Fatigue-PINN framework in the Unity engine, using Numpy and Keras Python wrappers for C#, the programming language for scripting in Unity.
  • Load our pre-trained models (i.e. initially train them in Python) in Unity using the same Python wrappers for C#.
  • Store the fatigued movement produced by our Fatigue-PINN, i.e. sequences of fatigued joint angles ($\acute{\hat{\mathbf{q}}}_{t}^{N}$), in a file format easily manipulated in C#, e.g. a TXT file.
The first two options involving the use of Python wrappers in C# result in frame-drops (by $\sim 10-15$ frames), since they require the initialization of the Python Engine as a background process, depleting the computational speed. Even though this may hinder the real-time execution of our Fatigue-PINN framework in Unity, our model can still be easily integrated into any motion synthesis pipeline operating on joint angles. The last option involves executing Fatigue-PINN as a standalone framework to produce fatigued motion sequences and subsequently loading the animation to a character.
The character utilized for showcasing our results in Unity is rigged. Character rigging is the 3D animation process where a set of manipulators and controls (e.g. joints, bones, etc.) are defined for a 3D virtual character. These controls constitute a ”rig” or a digital skeleton, which aids the animator to correctly render the animation on the character [80]. However, the joints in the skeleton of the character as well as their orientations, do not match with those of the model used for musculoskeletal modeling. Therefore, we manually altered the number of joints as well as their orientations in the local coordinate system (the one of the respective joint) to be aligned with the OpenSim model. This correspondence is essential for accurately endowing the animation on the character, since the fatigued joint angles produced by Fatigue-PINN are translated in the OpenSim’s local coordinate systems.
C# scripting is utilized to animate the virtual character. This process entails configuring the 3D local joint angles of the character with the values of the fatigued joint kinematics produced by our Fatigue-PINN framework.